### 24 Mar 2026 v1

@@ Abstract
Integrated opto-electronic devices have the potential to revolutionize information processing, with substantial increase in compute speed, seamless information transfer and reduction of energy consumption. A key missing unit for the successful implementation of compact functional devices are nanometer scale modular and tunable light sources. Monotonically grown semiconducting nanowire lasers (NWLs) fills this gap. However, NWLs operation improvement and optimization require the characterization of their near-field and its dynamics at the nanometer scale, which is hindered due to the light diffraction limit. Here we show how synchronous electron near-field and photon far-field time-resolved spectroscopies surpass this limitation and map a NWLs near-field with nanometer and sub-picoseconds temporal resolution. We quantitatively measured the evolution of the absolute number of stimulated photons in the NWL cavity, measuring that up to 4x105 are present simultaneously in the cavity. We mapped the lasing cavity mode’s near-field, showing that both whispering gallery and Fabry-Perot modes can participate in the lasing. Our results demonstrate how the near-field of a NWL under operation evolves in the sub-picoseconds and the nanometer scales. We anticipate that a direct observation of the near-field will help to elucidate the influence of materials heterogeneities (defects, chemical changes, contaminants, interface roughness, strain) in NWL operation.
@@
UTEM, NanoLaser, PINEM, light-matter interaction, out-of-equilibrium spectroscopySemiconductors nanowire lasers (NWLs) are one of the most promising new class of nanometer scale light sources, with sub-wavelength applications in bio-sensing and on-chip integrated photonics. Indeed, their monolithic growth insures both a high material quality and an economical viability [ellis_nanolasers_2024, behzadirad_scalable_2018, damilano_top-down_2019]. In addition, their confined cavity mode guarantee a low lasing threshold [gradecak_gan_2005, couteau_nanowire_2015] and ultrafast dynamics [roder_ultrafast_2015]. Following the initial experimental demonstration of a ZnO nanowire laser in 2001 [huang_room-temperature_2001], NWLs exploiting II-VI, III-V [johnson_single_2002, saxena_optically_2013], and various other materials [yang_room_2017] have been developed. The abundance of semiconductor gain materials enables a significant tunability for specific applications (wavelength, material compatibility, etc.). One of their most interesting characteristic is that the cavity properties are essentially determined by the geometry of the nanowire, due to the significant refractive index difference between vacuum and semiconductor medium [piprek_gan_2007]. This means that the lasing wavelength of a NWL is mainly dictated by the cavity length, a design parameter which can be easily tuned [piprek_gan_2007, zhuang_multicolor_2019]. However, despite the dramatic influence of these sub-wavelength parameters, the optical characteristics of nanolasers (e.g. lasing threshold, power density, mono-chromaticity, and lasing dynamics [roder_ultrafast_2015]) have up to now been investigated in the far-field region. There is thus a strong incentive to push the study of NWL in the near-field i.e. accessing the evanescent field with a subwavelength resolution. This was for example achieved on photonic crystal micro-lasers using scanning near-field optical microscopy (SNOM) [vo_near-field_2010], where an efficient mapping of the luminescence properties was reported. However, it was also shown that the near-field probe has a strong influence on the measured properties [le_gac_tuning_2009], hindering the measurement of intrinsic properties. In this article, we report the first synchronized measurement of far-field and spatially resolved sub-picoseconds temporal dynamics of the optical near-field of NWLs during lasing operation. To this aim, we use a combination of micro photo-luminescence spectroscopy (PL) and photon induced near-field electron microscopy (PINEM) [barwick_photon-induced_2009] implemented in a custom-built ultrafast transmission electron microscope (UTEM). In the following, we will demonstrate: (1) the quantitative measurement of the absolute number of stimulated photons inside the NWL cavity as a function of time with sub-picoseconds resolution, and (2) the mapping of the lasing cavity mode’s near-field with nanometer resolution.
@@ 1 PINEM on Nanowire laser
PINEM is a pump-probe technique where a focused pulsed electron beam probes the z component of the near-field induced by a pulsed laser on a nano-structure (see Fig. 1a). Before this work, in PINEM, the optical modes have always been resonantly excited by the pulsed laser and did not involved light absorption processes. Indeed here, the optical pumping produces electrons and holes relaxing and providing gain, leading to laser emission in a specific electromagnetic mode of the nanowire. When passing through the excited near-field of an object, the electron exchanges quanta of energy with the optical field, resulting in an electron energy spectrum with sidebands spaced by the probed photon energy (see Fig. 1d). The amplitude of the sideband - corresponding to the electron having exchanged with the field - is noted with . Remarkably, all the amplitudes are uniquely determined by a single quantity - the electron-light coupling constant - which can be expressed as [di_giulio_probing_2019]:
@@ 2.1 Temporal dynamics
We started by focusing the electron beam at a fixed position on the side of the NWL in an aloof geometry and recorded electron energy spectra for each delay between the pump laser pulse and the electron probe pulse (Fig. 1b), constructing what is called delay scans. These were recorded for a pumping below (P=0.25 mW) and above (P=0.66 mW) the lasing threshold (Fig. 2). For negative delays - corresponding to the probe impinging on the sample before the pump - the electron energy remains unchanged, and only the so-called zero-loss peak (ZLP) is observed, as expected.
@@ 2.2 Photon population
We attribute the first set of electron energy sidebands ( ps) to the pump laser light directly scattered at the surface of the nanowire, and the second set of sidebands to the stimulated near-field of the NWLs during lasing. We will now extract the electron-light coupling constant from the sidebands amplitude , starting from the first set of sidebands. Since they correspond to the interaction between an electron pulse (400 fs) which is longer than the light pulse (250 fs), the scattering can be considered incoherent (each electron of the wavepacket experience a different field). In this case, the sidebands are modeled by a Poisson distribution [park_photon-induced_2010]. The resulting fit of is shown on Fig. 3a for three different pump powers. This approximation does not apply to the second set of sidebands, where the optical near-field has a long dynamics compared to the electron wavepacket duration. The resulting slight increase in the coherence of the interaction was modeled as in [harvey_probing_2020] and the resulting is extracted for the same three different pump powers.
The coupling constant variations are attributed to the photon population dynamics in the cavity (equation (1)). To obtain a quantitative estimate of the photon population, we have measured the electron-photon coupling constant of the same NWL using EELS on a 200 keV monochromated STEM (NION Hermes). The EELS spectrum taken at the same position as the PINEM delay scan is shown on Fig. 3c, a peak at 3.3 eV is clearly visible just below the band gap (marked by a red arrow). More peaks are visible at lower energy corresponding to other cavity modes but not coupled to the bath of spontaneous photons and therefore unavailable for lasing. From this experiment, we have estimated a value of and deduced the photons population dynamic in the cavity for each pump powers (Fig 3-b).
@@ 2.3 Cavity mode imaging
So far, we have discussed the temporal dynamics of the photon inside the cavity, and will now turn to the spatial mapping of the near field of the lasing mode, with sub-wavelength resolution. Indeed, due to their large diameter (between 400 and 500 nm), the NWLs sustains two types of cavity excitation: Fabry-Perot modes (FPM) and whispering cavity modes (WGM) [coulon_gan_2012]. The far-field emission pattern is different for both kind of modes and are observed in PL experiments (see SI).
In order to characterize the near-field profiles of the bare cavity modes (i.e. in absence of external laser pumping), we performed an STEM-EELS hyperspectral imaging experiment on the same monochromated Nion Hermes machine - a technique which had already demonstrated its efficiency in probing such excitations [du_electron_2022, auad_unveiling_2022]. Fig. 4a shown a typical EELS study of these NWL. The spectrum taken at the side of the NWL exhibits clear peaks that are not visible at the top of the nanowire. We identified the WGM with the clear peaks present in the side while on the top of the NWL only a large number of FPM that almost seems continuous is visible [du_electron_2022]. Fitting the EELS spectrum around 3.3 eV we can separate the contribution to the WGM (amplitude of the peak) and to the Fabry Perrot (Amplitude of the continuous background). The different spatial variations of the cavity profile for FPW and WGM are clearly visible and shown on the Fig. 4a for E=3.3 eV.
@@ 3 Conclusion
It must be stressed that the spatial, temporal, and energy resolutions of the UTEM used in this studied are respectively limited to 10 nm, 300 fs, and 1.0 eV, and we therefore expect the new generation of high-resolution instruments [schroder_laser-driven_2025, castioni_nanosecond_2025] to be able to dive even deeper in the physics of nanolasers or into more complex geometries, such as SPASERs [ellis_nanolasers_2024]. For example, monochromated electron microscopes [castioni_nanosecond_2025] may be able to resolve challenges related to mode hopping, cavity mode interference, and inter-mode dynamics.
@@ Acknowledgements
The author thank Albert Polman and Nika Van Nielen for discussions and preliminary experiment. We also thank Mathieu Kociak for discussions and the all STEM group for their patience when the microscope was blocked at 200 keV. This work received support from the National Agency for Research under the program of future investment TEMPOS-CHROMATEM (ANR-10- EQPX-50). The authors acknowledge financial support from the CNRS-CEA “METSA” French network (FR CNRS 3507) on the platform LPS-STEM. The author acknowledge financial support of the French Agence Nationale de la Recherche (ANR), under the grant agreement ANR-23-CE09-0018 - LUTEM.

### 26 Mar 2026 v2

@@ Abstract
Integrated opto-electronic devices have the potential to revolutionize information processing, with substantial increase in computing speed, seamless information transfer and reduction of energy consumption. A key missing unit for the successful implementation of compact functional devices are nanometer scale modular and tunable light sources. Monotonically grown semiconducting nanowire lasers (NWLs) fill this gap. However, NWLs operation improvement and optimization require the characterization of their near-field and its dynamics at the nanometer scale, which is hindered due to the light diffraction limit. Here we show how synchronous electron near-field and photon far-field time-resolved spectroscopies surpass this limitation and map a NWLs near-field with nanometer and sub-picoseconds temporal resolution. We quantitatively measured the evolution of the absolute number of stimulated photons in the NWL cavity, measuring that up to 4x105 are present simultaneously in the cavity. We mapped the lasing cavity mode’s near-field, showing that both whispering gallery and Fabry-Perot modes can participate in the lasing. Our results demonstrate how the near-field of a NWL under operation evolves in the sub-picoseconds and the nanometer scales. We anticipate that a direct observation of the near-field will help to elucidate the influence of materials heterogeneities (defects, chemical changes, contaminants, interface roughness, strain) in NWL operation.
@@
UTEM, NanoLaser, PINEM, light-matter interaction, out-of-equilibrium spectroscopySemiconductors nanowire lasers (NWLs) are one of the most promising new class of nanometer scale light sources, with sub-wavelength applications in bio-sensing and on-chip integrated photonics. Indeed, their monolithic growth insures both a high material quality and an economical viability [ellis_nanolasers_2024, behzadirad_scalable_2018, damilano_top-down_2019]. In addition, their confined cavity mode guarantees a low lasing threshold [gradecak_gan_2005, couteau_nanowire_2015] and ultrafast dynamics [roder_ultrafast_2015]. Following the initial experimental demonstration of a ZnO nanowire laser in 2001 [huang_room-temperature_2001], NWLs exploiting II-VI, III-V [johnson_single_2002, saxena_optically_2013], and various other materials [yang_room_2017] have been developed. The abundance of semiconductor gain materials enables a significant tunability for specific applications (wavelength, material compatibility, etc.). One of their most interesting characteristic is that the cavity properties are essentially determined by the geometry of the nanowire, due to the significant refractive index difference between vacuum and semiconductor medium [piprek_gan_2007]. This means that the lasing wavelength of a NWL is mainly dictated by the cavity length, a design parameter which can be easily tuned [piprek_gan_2007, zhuang_multicolor_2019]. However, despite the dramatic influence of these sub-wavelength parameters, the optical characteristics of nanolasers (e.g. lasing threshold, power density, mono-chromaticity, and lasing dynamics [roder_ultrafast_2015]) have up to now only been investigated in the far-field region. There is thus a strong incentive to push the study of NWL in the near-field i.e. accessing the evanescent field with a subwavelength resolution. This was for example achieved on photonic crystal micro-lasers using scanning near-field optical microscopy (SNOM) [vo_near-field_2010], where an efficient mapping of the luminescence properties was reported. However, it was also shown that the near-field probe has a strong influence on the measured properties [le_gac_tuning_2009], hindering the measurement of intrinsic properties. In this article, we report the first synchronized measurement of far-field and spatially resolved sub-picoseconds temporal dynamics of the optical near-field of NWLs during lasing operation. To this aim, we use a combination of micro photo-luminescence spectroscopy (PL) and photon induced near-field electron microscopy (PINEM) [barwick_photon-induced_2009] implemented in a custom-built ultrafast transmission electron microscope (UTEM). In the following, we will demonstrate: (1) the quantitative measurement of the absolute number of stimulated photons inside the NWL cavity as a function of time with sub-picoseconds resolution, and (2) the mapping of the lasing cavity mode’s near-field with nanometer resolution.
@@ 1 PINEM on Nanowire laser
PINEM is a pump-probe technique where a focused pulsed electron beam probes the z component of the near-field induced by a pulsed laser on a nano-structure (see Fig. 1a). Before this work, in PINEM, the optical modes have always been resonantly excited by the pulsed laser and did not involved light absorption processes. Indeed here, the optical pumping produces electrons and holes relaxing and providing gain, leading to laser emission in a specific electromagnetic mode of the nanowire. When passing through the excited near-field of an object, the electron exchanges quanta of energy with the optical field, resulting in an electron energy spectrum with sidebands spaced by the probed photon energy (see Fig. 1d). The amplitude of the sideband - corresponding to the electrons having exchanged with the field - is noted with . Remarkably, all the amplitudes are uniquely determined by a single quantity - the electron-light coupling constant - which can be expressed as [di_giulio_probing_2019]:
@@ 2.1 Temporal dynamics
We started by focusing the electron beam at a fixed position on the side of the NWL in an aloof geometry and recorded electron energy spectra for each delay between the pump laser pulse and the electron probe pulse (Fig. 1b), constructing so-called delay scans. These were recorded for a pumping below (P=0.25 mW) and above (P=0.66 mW) the lasing threshold (Fig. 2). For negative delays - corresponding to the probe impinging on the sample before the pump - the electron energy remains unchanged, and only the so-called zero-loss peak (ZLP) is observed, as expected.
@@ 2.2 Photon population
We attribute the first set of electron energy sidebands ( ps) to the pump laser light directly scattered at the surface of the nanowire, and the second set of sidebands to the stimulated near-field of the NWLs during lasing. We will now extract the electron-light coupling constant from the sidebands amplitude , starting from the first set of sidebands. Since they correspond to the interaction between an electron pulse (400 fs) which is longer than the light pulse (250 fs), the scattering can be considered incoherent (each electron of the wavepacket experiences a different field). In this case, the sidebands are modeled by a Poisson distribution [park_photon-induced_2010]. The resulting fit of is shown on Fig. 3a for three different pump powers. This approximation does not apply to the second set of sidebands, where the optical near-field has a long dynamics compared to the electron wavepacket duration. The resulting slight increase in the coherence of the interaction was modeled as in [harvey_probing_2020] and the resulting is extracted for the same three different pump powers.
The variations of the coupling constant are attributed to the photon population dynamics in the cavity (equation (1)). To obtain a quantitative estimate of the photon population, we have measured the electron-photon coupling constant of the same NWL using EELS on a 200 keV monochromated STEM (NION Hermes). The EELS spectrum taken at the same position as the PINEM delay scan is shown on Fig. 3c, a peak at 3.3 eV is clearly visible just below the band gap (marked by a red arrow). More peaks are visible at lower energy corresponding to other cavity modes but not coupled to the bath of spontaneous photons and therefore unavailable for lasing. From this experiment, we have estimated a value of and deduced the photons population dynamic in the cavity for each pump powers (Fig 3-b).
@@ 2.3 Cavity mode imaging
So far, we have discussed the temporal dynamics of the photon inside the cavity, and will now turn to the spatial mapping of the near field of the lasing mode with sub-wavelength resolution. Indeed, due to their large diameters (between 400 and 500 nm), the NWLs sustain two types of cavity excitation: Fabry-Perot modes (FPM) and whispering cavity modes (WGM) [coulon_gan_2012]. The far-field emission pattern is different for both kind of modes and are observed in PL experiments (see SI).
In order to characterize the near-field profiles of the bare cavity modes (i.e. in absence of external laser pumping), we performed an STEM-EELS hyperspectral imaging experiment on the same monochromated Nion Hermes machine - a technique which had already demonstrated its efficiency in probing such excitations [du_electron_2022, auad_unveiling_2022]. Fig. 4a shows a typical EELS study of these NWL. The spectrum taken at the side of the NWL exhibits clear peaks that are not visible at the top of the nanowire. We identified the WGM with the clear peaks present in the side while on the top of the NWL only a large number of FPM that almost seems continuous is visible [du_electron_2022]. Fitting the EELS spectrum around 3.3 eV we can separate the contribution of the WGM (amplitude of the peak) and of the Fabry Perrot (Amplitude of the continuous background). The different spatial variations of the cavity profile for FPW and WGM are clearly visible and shown on the Fig. 4a for E=3.3 eV.
@@ 3 Conclusion
It must be stressed that the spatial, temporal, and energy resolutions of the UTEM used in this study are respectively limited to 10 nm, 300 fs, and 1.0 eV, and we therefore expect the new generation of high-resolution instruments [schroder_laser-driven_2025, castioni_nanosecond_2025] to be able to dive even deeper in the physics of nanolasers or into more complex geometries, such as SPASERs [ellis_nanolasers_2024]. For example, monochromated electron microscopes [castioni_nanosecond_2025] may be able to resolve challenges related to mode hopping, cavity mode interference, and inter-mode dynamics.
@@ Acknowledgements
The author thank Albert Polman and Nika Van Nielen for discussions and preliminary experiments. We also thank Mathieu Kociak for discussions and the all STEM group for their patience when the microscope was blocked at 200 keV. This work received support from the National Agency for Research under the program of future investment TEMPOS-CHROMATEM (ANR-10- EQPX-50). The authors acknowledge financial support from the CNRS-CEA “METSA” French network (FR CNRS 3507) on the platform LPS-STEM. The author acknowledge financial support of the French Agence Nationale de la Recherche (ANR), under the grant agreement ANR-23-CE09-0018 - LUTEM.